OpenStudy (anonymous):

Complete the square for the binomial:x square+4x

OpenStudy (anonymous):

what is 1/2 of 4 = 2 now square that = 4

OpenStudy (anonymous):

simple: x² + 4x + 4 = (x + 2)²

OpenStudy (anonymous):

resolves?

OpenStudy (anonymous):

\[x^2+4x+(\frac{4}{2})^2-(\frac{4}{2})^2=x^2+4x+4-4=(x+2)^2-4\]

OpenStudy (anonymous):

Don't forget to subtract the same value you added to complete the square, otherwise you won't have an equivalent expression

OpenStudy (anonymous):

I think that is where I keep messin up

OpenStudy (anonymous):

Fonez I have one more